The difference of the squares of two consecutive positive even integers is 20. What is the average of the two numbers?

Answer by mananth(16946)


Let x and x+2 be the numbers
(x+2)^2-x^2=20
x^2+4x+4 -x^2=20
4x=16
x=4
x+2=6
4 & 6 are the numbers.
.
4+6 / 2 = 5 is the average